The Hill Company reported the following results:

Year 3 Year 2 Year 1
Income Statement
Revenue 10,972 11,598 10,470
Cost of Goods Sold 8,942 8,767 7,901
Selling, General & Admin. Expenses 2,470 2,611 2,479
Interest expense 76 80 28
Net Income (516) 140 62
Balance Sheet
Assets
Cash 1,354 1,316 1,880
Prepaid expenses 202 522 125
Accounts receivable 375 250 231
Inventory 745 698 455
Property & equipment (net) 20,464 18,810 17,727
Total Assets 23,140 21,596 20,418
Liabilities
Accounts payable 2,824 743 678
Unredeemed gift cards 410 850 636
Notes Payable 15,457 18,048 17,024
Stockholders' Equity
Common Stock 985 545 815
Retained Earnings 3,464 1,410 1,265
Total Liabilities & Equity 23,140 21,596 20,418

1. What is the company's gross profit ratio for Year 3? Convert your final answer to a percentage, round to one decimal place and enter without the "%" sign (e.g. a final answer of 0.105678 would be entered as 10.6).

2. What is the company's return on investment ratio for Year 1? Convert your final answer to a percentage, round to one decimal place and enter without the "%" sign (e.g. a final answer of 0.105678 would be entered as 10.6).

3. Based on the three years of data, the company's return on investment ratio has

a) improved.

b) worsened.

c) stayed the same.
Answer #1

Answer:

Gross Profit % = Gross Profit / Sales

Year 3 = ( 10972 - 8942 ) / 10972 = 18.5%

Year 2 = ( 11598 - 8767 ) / 11598 = 24.41%

Year 1 = ( 10470 - 7901 ) / 10470 = 24.54%

Return on Investment = Net Income / (Long Term Liabilities + Stockholders Equity

Where Note Payable is Current Liability

Year 3 = -516 / (985 + 3464) = -11.6%

Year 2 = 140 / ( 545 + 1410 ) = 7.16%

Year 1 = 62 / ( 815 + 1265 ) = 2.98%

The ratio improved in Year 2 but failed miserably or worsened in Year 3

Where Note Payable is Long Term Liability

Year 3 = -516 / (985 + 3464 + 15457 ) = -2.6%

Year 2 = 140 / ( 545 + 1410 + 18048) = 0.699%

Year 1 = 62 / ( 815 + 1265 + 17024 ) = 0.324%